These I am completely lost on - could someone please help with these two? The rest will fall into place if I have something to go on.
Simplify the rational expressions
-x^2 = 10x - 21
________________
-3x^2 + 23x - 14
and
28x^2 - 29x + 6
________________
-28x^2 + 50x - 12
again, thank you very much!

Answer by checkley71(8403)

-X^2=10X-21
-X^2-10X+21=0
X^2+10X-21=0
USING THE QUADRATIC EQUATION WE GET
X=(-10+-SQRT(100-4*-21))/2
X=(-10+-SQRT(100+84))/2
X=(-10+-SQRT184)/2
X=(-10+-13.56)/2
x=(-10+13.56)/2
x=3.56/2
x=1.78
and
x=(-10-13.56)/2
x=-23.56/2
x=-11.78